OpenStudy (anonymous):

lim as x-->0 of xcscx Can someone show steps please.

OpenStudy (agreene):

\[\lim_{x\rightarrow 0}[ x\csc(x)] =\lim_{x \rightarrow 0}[\frac{x}{\sin(x)}]\] the second is indeterminate, 0/0. So, apply L'Hopitals: \[\lim_{x\rightarrow 0} \frac{\frac{d}{dx}x}{\frac{d}{dx}\sin(x)}\implies \lim_{x \rightarrow 0} \frac{1}{\cos(x)}=\frac{1}{ \lim_{x \rightarrow 0}\cos(x)}=\frac{1}{1}=1\]

OpenStudy (anonymous):

Thank you very much. I missed an important thing.

OpenStudy (agreene):

np

OpenStudy (anonymous):

Can i ask you one more question

OpenStudy (agreene):

sure

OpenStudy (anonymous):

wait, never mind i got it! Thanks

OpenStudy (agreene):

lol yw

OpenStudy (zarkon):

if you know \[\lim_{x \rightarrow 0}\frac{\sin(x)}{x}=1\] then \[\lim_{x\rightarrow 0} x\csc(x) =\lim_{x \rightarrow 0}\frac{x}{\sin(x)}=\lim_{x \rightarrow 0}\frac{1}{\frac{\sin(x)}{x}}=\frac{1}{1}=1\]

OpenStudy (anonymous):

is this true for tanx/x=1?

OpenStudy (zarkon):

\[\frac{\tan(x)}{x}=\frac{1}{\cos(x)}\frac{\sin(x)}{x}\]

OpenStudy (zarkon):

\[\to \frac{1}{1}\cdot1=1\]

OpenStudy (zarkon):

as \(x\to 0\)

OpenStudy (anonymous):

hey @Zarkon can i ask you one more thing?

OpenStudy (zarkon):

?

OpenStudy (anonymous):

Can you explain this question: